OpenStudy (anonymous):

Which fraction expresses in lowest terms 5/12?

OpenStudy (anonymous):

I think your fraction is simplified to it's lowest term. :)

OpenStudy (anonymous):

A. 2/3

OpenStudy (anonymous):

B. 10/18

OpenStudy (anonymous):

It's already in lowest terms.

OpenStudy (cwrw238):

yes - there are no common factors for 5 and 12

OpenStudy (anonymous):

C. 20/54

OpenStudy (anonymous):

OH i messed up the question

OpenStudy (anonymous):

Lel

OpenStudy (anonymous):

Oh do you mean which fraction when simplified to it lowest terms equals 5/12?

OpenStudy (anonymous):

Which fraction is equivalent to

OpenStudy (anonymous):

Ohhhh

OpenStudy (cwrw238):

try writing them in prime factor form

OpenStudy (anonymous):

5/18

OpenStudy (cwrw238):

20 = 2*2*5 54 = ?

OpenStudy (cwrw238):

can you factor 54?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

What do you mean by that?

OpenStudy (cwrw238):

stsrt first with 2 as 54 is even then try 3,5 etc

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

54 / 2 = 27

OpenStudy (cwrw238):

27 / 3 = ? then keep on until the last quotient is 1

OpenStudy (anonymous):

9

OpenStudy (cwrw238):

right i'll write it out